OpenStudy (anonymous):

Mike went to a football game with his family. They bought 5 boxes of popcorn, 3 programs and spent $41. Sean's family went to the same game, bought 3 boxes of popcorn, 2 programs and spent $26. How much does each program cost at the game?

OpenStudy (anonymous):

who was playing?

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

ok \[5x+3y=41\] \[3x+2y=26\]

OpenStudy (anonymous):

let the cost be x and y for popcorn and program 5x + 3y=41 2x + 2y = 26 Now solve them

OpenStudy (anonymous):

but 3 dosnt gointo 26

OpenStudy (anonymous):

\[10x+6y=82\] \[-9x-6y=-78\] add to get \[x=4\]

OpenStudy (anonymous):

ah sorry 3x+2y = 26

OpenStudy (anonymous):

y=7

OpenStudy (anonymous):

ok i used your formula i agree

OpenStudy (anonymous):

and since \[x=4\] we know \[5\times 4+3y=41\] \[20+3y=41\] \[3y=21\] \[y=7\]

OpenStudy (anonymous):

popcorn is $3, programs are $7 and it must be a high school game because otherwise these prices would be doubled at least

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new equation after the expression equivalent to x from the first equation is substituted into the second equation. x + y = 2 3x + 2y = 9

OpenStudy (anonymous):

does anyone want to give this a shot